# Towards a Theoretical Understanding of Memorization in Diffusion Models

### v1

@@ 5.3 Main Results
The "Random" baseline generates images directly using the target unconditional DPM, as described in (Carlini et al., 2023). We average the results across various values of $\lambda$ (defined in Eq. (151)), ranging from 5 to 9, with a detailed analysis provided in Section 5.3. It is important to note that $\lambda=0$ corresponds to the "Random" baseline.
@@
Here, we test the sensitivity of SIDE to its hyper-parameter $\lambda$. We generate 50,000 images for each integer value of $\lambda$ within the range of [0, 50]. As shown in Figure 4, the memorization score increases at first, reaching its highest, then decreases as $\lambda$ increases. This can be understood from sampling SDE Eq. (151). Starting from 0, the diffusion models are unconditional. As $\lambda$ increases, the diffusion models become conditional, and according to Theorem 2, the memorization effect will be triggered. However, when $\lambda$ is too large, the generated images will overfit the classifier’s decision boundaries, leading to a low diversity and ignoring the data distribution. Consequently, the memorization score decreases.
@@ A.4 Proof for theorem 2
According to the Eq. (8), the nuclear norm of the two covariance matrices differs. Specifically:
@@ Appendix B Explanation of the $N_{G}$’s importance in fair comparison
The importance of $N_{G}$ lies in its impact on the non-linear expectation of uniquely memorized images. Comparing UMS across different $N_{G}$ values is flawed because varying $N_{G}$ leads to different outcomes, underscoring the need for consistent $N_{G}$ values to ensure fair comparisons. In Carlini et al. (2023), the significance of $N_{G}$ was overlooked, as they only reported the number of uniquely extracted images. Our theoretical analysis accurately aligns with the behavior observed in the experimental data in Fig 5.
@@ Appendix E Pseudocode for SIDE method
20:         Sample $\bm{x}_{0}$ conditioned on $y=c$ using Eq. (151)

### v2

@@ 5.3 Main Results
The "Random" baseline generates images directly using the target unconditional DPM, as described in carlini2023extracting. We average the results across various values of $\lambda$ (defined in equation (151)), ranging from 5 to 9, with a detailed analysis provided in Section 5.3. It is important to note that $\lambda=0$ corresponds to the "Random" baseline.
@@
Here, we test the sensitivity of SIDE to its hyper-parameter $\lambda$. We generate 50,000 images for each integer value of $\lambda$ within the range of [0, 50]. As shown in Figure 4, the memorization score increases at first, reaching its highest, then decreases as $\lambda$ increases. This can be understood from sampling SDE Equation (151). Starting from 0, the diffusion models are unconditional. As $\lambda$ increases, the diffusion models become conditional, and according to Theorem 2, the memorization effect will be triggered. However, when $\lambda$ is too large, the generated images will overfit the classifier’s decision boundaries, leading to a low diversity and ignoring the data distribution. Consequently, the memorization score decreases.
@@ 6 Conclusion
Broader Impacts
Our work proposes a training data extraction method targeted at diffusion models and the effectiveness of our proposed method has far exceeded random methods. This may be exploited by potential malicious users to attack existing generative diffusion models. By effectively avoiding the introduction of additional or unnecessary informative label information into the training process, the risk of a diffusion model being attacked by data extraction can be greatly reduced. Therefore, the contribution of our work to the community outweighs the risks it poses, especially considering that the scale of our extraction method has not reached the same scale of state-of-the-art diffusion models. Moreover, the theoretical and experimental understandings established in this work can help the community build a deeper understanding of the memorization mechanism of diffusion models, and inspire better defense methods.
[…]
@@ A.4 Proof for theorem 2
According to the equation (8), the nuclear norm of the two covariance matrices differs. Specifically:
@@ Appendix B Explanation of the $N_{G}$’s importance in fair comparison
The importance of $N_{G}$ lies in its impact on the non-linear expectation of uniquely memorized images. Comparing UMS across different $N_{G}$ values is flawed because varying $N_{G}$ leads to different outcomes, underscoring the need for consistent $N_{G}$ values to ensure fair comparisons. In carlini2023extracting, the significance of $N_{G}$ was overlooked, as they only reported the number of uniquely extracted images. Our theoretical analysis accurately aligns with the behavior observed in the experimental data in Fig 5.
@@ Appendix E Pseudocode for SIDE method
20:         Sample $\bm{x}_{0}$ conditioned on $y=c$ using equation (151)